I am having trouble with the following problem:
y=4x+11
3x-2y=-7

Answer by AnlytcPhil(1806)

 y = 4x + 11
3x - 2y = -7

The first equation is already solved for y, so replace 
the y in the second equation by what y equals, namely,
(4x+11)

      3x - 2y = -7
3x - 2(4x+11) = -7
 3x - 8x - 22 = -7
     -5x - 22 = -7
          -5x = 15
            x = -3

Now to find y, replace x by -3 in the first equation:

   3x - 2y = -7
3(-3) - 2y = -7
   -9 - 2y = -7
       -2y = 2
         y = -1

So the solution is (x, y) = (-3, -1)

Edwin
AnlytcPhil@aol.com